OpenStudy (anonymous):

Solve by the elimination method. 0.3x-0.2y=4 0.2x+0.3y= -37/17

OpenStudy (saifoo.khan):

u did this b4 too.

OpenStudy (saifoo.khan):

x = 32.7 y = -29

OpenStudy (anonymous):

(32.7,-29) ?

OpenStudy (saifoo.khan):

Yes!

OpenStudy (anonymous):

ok. I did do a similar one before. that answer was a fraction though/

OpenStudy (amistre64):

id rather ti83 it :)

OpenStudy (anonymous):

huh?

OpenStudy (saifoo.khan):

o.O

OpenStudy (saifoo.khan):

amistre, is busy incresing his xp! ;) jk

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

:) gotta gain up lost points from the outtage lol

OpenStudy (amistre64):

the ti83 gives me decimals; and this thinl looks to turn mean in the end

OpenStudy (amistre64):

it close to (6,-11) :) if that helps

OpenStudy (amistre64):

i fthis is on a program; they give any options?

OpenStudy (anonymous):

lol so the answer is decimals? I juist want to make sure it accepts the answer

OpenStudy (amistre64):

the answers are fractions for exact form; but my calc only gives me the decimals

OpenStudy (amistre64):

*10; 0.3x-0.2y=4 *10; 0.2x+0.3y= -37/17 3x-2y=40 2x+3y= -370/17 ; *17 3x-2y=40 34x+51y= -370 this looks better and is prolly more doable

OpenStudy (anonymous):

it says simplify your answer. Type an ordered pair. Use integers or fractions for any numbers in the expression.

OpenStudy (amistre64):

sounds like mathlab stuff

OpenStudy (anonymous):

LOL

OpenStudy (saifoo.khan):

haha.

OpenStudy (amistre64):

3x- 2y= 40 ; *34 34x+51y= -370; * -3 102x - 68y = 1360 -102x -153y = 1110 -------------------- 221y = 2470 y = 2470/221 cant tell if it simplifies yet :)

OpenStudy (amistre64):

y = 190/17

OpenStudy (anonymous):

wow...I am going to study over this while you figure it out. There is still x to solve right?

OpenStudy (amistre64):

3x - 2y = 40 ; *51 34x +51y = -370; * 2 153x -102y = 2040 68x +102y= -740 ------------------- 221x = 1300 x = 1300/221 ; /13 top to bottom x = 100/17

OpenStudy (amistre64):

i think i tripped over myself; gotta dbl chk

OpenStudy (amistre64):

i cant get them to match up :/

OpenStudy (amistre64):

i see it; i dropped a negative sign lol

OpenStudy (amistre64):

that got it lol

OpenStudy (amistre64):

x=100/17 y=-190/17

OpenStudy (anonymous):

thank you.

OpenStudy (amistre64):

yep; hunting down that lost - was a bummer :) everything else was good

OpenStudy (anonymous):

I was lost. LOL